OpenStudy (anonymous):

Gregory teaches martial arts. He charges a one-time processing fee of $5.00 and the cost of the classes is shown below. Let x represent the number of classes and y represent the cost of classes. Based on this information, what will it cost to take 10 classes (including processing fee)? Cost of classes (not including processing fee) Number of Classes, x 1 2 3 4 Cost of Classes, y $15.00 $27.00 $39.00 $51.00

OpenStudy (anonymous):

Medals for all

OpenStudy (aravindg):

:) mmmeeeeeee

OpenStudy (anonymous):

lol, You have to help me with my work, & every time you help i will give you a medal on each question

OpenStudy (anonymous):

$5 + $15 + 9*$12 = 20 + 108 = $128

OpenStudy (anonymous):

Actually, $5 less without including processing fee. So, $123.

OpenStudy (anonymous):

And, $128 including the processing fee of $5.